A bag contains {6} red marbles, {5} yellow marbles, and {7} green marbles. How many additional red marbles must be added to the {18} marbles alredy in the bag so that the probability of randomly drawing a red marble is {3/5}?
Answer by checkley77(12569)

6+5+7=18 total marbles.
3/5=(6+x)/(18+x)
5(6+x)=3(18+x)
30+5x=54+3x
5x-3x=54-30
2x=24
x=24/2
x=12 red marbles need to be added.
Proof:
3/5=(6+12)/(18+12)
3/5=18/30
3/5=3/5